OpenStudy (anonymous):

what is the anti derivative of x/(x^2+1) and why?

hartnn (hartnn):

try u= x^2+1 du= ..?

OpenStudy (anonymous):

du= 2x

hartnn (hartnn):

only 2x ? or du = 2x dx ?

OpenStudy (anonymous):

yes 2x dx

hartnn (hartnn):

so, x dx = du/2 so whats your new integral in terms of 'u' only ?

OpenStudy (anonymous):

2x/2 dx

OpenStudy (anonymous):

or 2x dx/2..

hartnn (hartnn):

what ?

hartnn (hartnn):

sorry i ddn't get what doubt you have ..

OpenStudy (anonymous):

so in respect to u it would be 2x/2 then

hartnn (hartnn):

i am not sure on how you get 2x/2 let me show you steps

hartnn (hartnn):

u = x^2+1 du= 2x dx xdx = du/2 which is your numerator, denominator = x^2+1 = u so your new integral is \(\int \dfrac{du}{2u}\) got this ?

OpenStudy (anonymous):

ok yes

hartnn (hartnn):

can you integrate that ?

hartnn (hartnn):

\(\int (1/x)dx = \ln |x|+c \)

OpenStudy (anonymous):

the answer on the example is \[\lim_{b \rightarrow \infty} [\frac{ 1 }{ 2}\ln(x ^{2}+1)]_{1}^{b}\]

hartnn (hartnn):

ok

hartnn (hartnn):

can you integrate \(\int \dfrac{du}{2u}\) ?

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \frac{ 1 }{ 2 }*\frac{ 1 }{ x^2+1 }= 1/2 \ln(x^2+1)\]

hartnn (hartnn):

\(\int du/2u = 1/2 \ln |u|+c\) but u = x^2+1 so, \(= 1/2 \ln |x^2+1| +c\)

hartnn (hartnn):

and if you are given limits from 1 to infinity you get \(\lim_{b \rightarrow \infty} [\frac{ 1 }{ 2}\ln(x ^{2}+1)]_{1}^{b}\)

OpenStudy (anonymous):

thank you, btw what is the rule that you used earlier called? the du/u stuff?

hartnn (hartnn):

its not called anything, its just a formula...

OpenStudy (anonymous):

gotcha